# How Far Are We on the Decision-Making of LLMs? Evaluating LLMs’ Gaming Ability in Multi-Agent Environments

### v1

@@ Abstract
Decision-making, a complicated task requiring various types of abilities, presents an excellent framework for assessing Large Language Models (LLMs). Our research investigates LLMs’ decision-making capabilities through the lens of a well-established field, Game Theory. We focus specifically on games that support the participation of more than two agents simultaneously. Subsequently, we introduce our framework, $\gamma$-Bench, including eight classical multi-agent games. We design a scoring scheme to assess a model’s performance in these games quantitatively. Through $\gamma$-Bench, we investigate LLMs’ robustness, generalizability, and enhancement strategies. Results reveal that while GPT-3.5 shows satisfying robustness, its generalizability is relatively limited. However, its performance can be improved through approaches such as Chain-of-Thought. Additionally, we conduct evaluations across various LLMs and find that GPT-4 outperforms other models on $\gamma$-Bench, achieving a score of 72.5. Moreover, the increasingly higher scores across the three iterations of GPT-3.5 (0613, 1106, 0125) demonstrate marked advancements in the model’s intelligence with each update. The code and experimental results are made publicly available via https://github.com/CUHK-ARISE/GAMABench.
@@ 1 Introduction
Public Goods Game $82.3_{\pm 1.7}$ $56.5_{\pm 12.6}$ $61.2_{\pm 8.1}$ $10.8_{\pm 1.8}$ $31.5_{\pm 7.6}$
Diner’s Dilemma $33.0_{\pm 4.9}$ $98.6_{\pm 1.3}$ $97.2_{\pm 2.8}$ $99.1_{\pm 0.7}$ $96.9_{\pm 1.5}$
Sealed-Bid Auction $89.8_{\pm 0.4}$ $90.3_{\pm 1.5}$ $86.7_{\pm 1.6}$ $85.6_{\pm 2.4}$ $76.8_{\pm 4.3}$
@@ 2.1 Game Theory
Nash Equilibrium
Theorem 2.1 (Nash’s Existence Theorem)
Human Behaviors
@@
Duan et al. (2024) gpt-3.5-turbo, gpt-4, 0.2 Ten Games
@@
This Study gpt-3.5-turbo, gpt-4 0$\sim$1 Eight Games
@@ 3.2 Betraying Games
The Sealed-Bid Auction (SBA) involves players submitting their bids confidentially and simultaneously, different from the auctions where bids are made openly in a sequential manner. We consider two variants of SBA: the First-Price Sealed-Bid Auction (FPSBA) and the Second-Price Sealed-Bid Auction (SPSBA). In FPSBA, also known as the Blind Auction, if all players bid their true valuation $v_{i}$ of the item, the winner achieves a net gain of $v_{i}-v_{i}=0$ while others also gain nothing McAfee & McMillan (1987). Moreover, the highest bidder will discover that to win the auction, it is sufficient to bid marginally above the second-highest bid. Driven by these two factors, FPSBA is often deemed inefficient in practical scenarios, as bidders are inclined to submit bids significantly lower than their actual valuation, resulting in suboptimal social welfare. In contrast, SPSBA, commonly called the Vickrey auction, requires the winner to pay the second-highest bid, encouraging truthful bidding by all players Vickrey (1961). It can be proven that bidding true valuations in SPSBA represents an NE. This auction evaluates agent performance in imperfect information games, where agents lack knowledge of other players’ valuations.
@@ 4 Vanilla Experiments
Key Findings: The model’s decisions are mainly influenced by the outcomes of the preceding round rather than deriving from the reasoning of the optimal strategy. Although initially demonstrating suboptimal performance, the model can learn from historical data and enhance its performance over time. The model demonstrates the ability to engage in spontaneous cooperation, leading to increased social welfare beyond mere self-interest, without the necessity for explicit communication. The model shows limitations in sequential games characterized by more complicated rules. The aggregate score of the model on $\gamma$-Bench is $68.8$.
@@
System
You are participating in a game played by $N$ players over $K$ rounds.
@@
User
Game Results for Round $I$:
@@
Assistant
{“chosen_number”: “$C_{IJ}$”}
User
[Congratulation you won]/[Unfortunately you lost].
@@
User
Now round $I$ starts.
@@ 4.2 Betraying Games
The vanilla setting for this game is $R=2$. Each player has $T=20$ to contribute in each round. Fig. 5 shows the contributed tokens by each agent and their corresponding gains per round. The observations reveal the following: (1) Despite an investment return of $-80\%$, agents display a pattern of alternating between free-riding and contributing all their tokens. (2) As the rounds progress, there is an evident increase in the number of tokens contributed to the public pot, leading to an overall enhancement in social welfare gains. These findings suggest that the LLM exhibits cooperative behavior, prioritizing collective benefits over individual self-interest. Since we expect the model to maximize social welfare beyond mere self-interests, the optimal strategy should be contributing $T$. The raw score in this game is then given by $S_{4}=\frac{1}{NK}\sum_{ij}C_{ij}$, where $C_{ij}=1$ is the proposed contribution amount of player $i$ in round $j$. The model scores $58.8$ on this game.
@@
The vanilla setting for this game is $P_{h}=20$, $P_{l}=10$, $U_{h}=20$, $U_{l}=15$. We show the probability of agents choosing the costly dish, their resulting utilities, and the average bill in Figure 6. Analysis of the figure reveals the following insights: (1) Contrary to the NE predictions for this game, agents predominantly prefer the cheap dish, which maximizes total social welfare. (2) Remarkably, a deviation from cooperative behavior is observed wherein one agent consistently chooses to betray others, thereby securing a higher utility. This pattern of betrayal by this agent persists across subsequent rounds. Since we expect the model to maximize social welfare beyond mere self-interests, the optimal strategy should be choosing the cheap dish. The raw score in this game is then given by $S_{5}=\frac{1}{NK}\sum_{ij}D_{ij}$, where $D_{ij}=1$ when player $i$ chose the cheap dish in round $j$ and $D_{ij}=0$ when player $i$ chose the costly dish. The model scores $96.0$ on this game.
@@
For the vanilla setting in this game, we randomly assign valuations to each agent in each round, ranging from 00 to $200$. We fix the seed for random number generation to ensure fair comparisons across various settings and models. We evaluate LLMs’ performance under both First-Price and Second-Price settings. Fig. 7 depicts the subtraction between valuations and bids and bid amounts of each agent. Our key findings include: (1) As introduced in §3.2, we note that agents generally submit bids that are lower than their valuations in the First-Price auction, a tendency indicated by the positive discrepancies between valuations and bids depicted in Fig. 7(a). (2) Though the NE suggests that everyone bids the amount of their valuation in the Second-Price setting, we find a propensity for bidding below valuation levels, as demonstrated in Fig. 7(c). Since the optimal strategy is to bid their true valuations, the raw score [^4] For simplicity, we evaluate only the Second-Price setting. in this game is given by $S_{6}=\frac{1}{NK}\sum_{ij}(v_{ij}-b_{ij})$, where $v_{ij}$ and $b_{ij}$ are player $i$’s valuation and bid in round $j$, respectively. The model scores $88.3$ on this game.
@@ 5.1 RQ1: Robustness
We also investigate the impact of prompt phrasing on model performance. Leveraging GPT-4, we rewrite our default prompt templates described in §A, generating four additional versions. We perform a manual checking process on the generated versions to ensure GPT-4’s adherence to game rules without altering critical data. The prompt templates can be found in §B in the appendix. We plot the results of using these templates in Fig. 11 and record the quantitative scores in Table 7 in the Appendix. Notably, our findings indicate that prompt wording can significantly affect performance, as evidenced by the declines observed in Fig. 11(1), (5), and (6).
@@ 5.2 RQ2: Reasoning Strategies
According to Kojima et al. (2022), introducing a preliminary phrase, “Let’s think step by step,” encourages the model to sequentially analyze and explain its reasoning before presenting its conclusion. This approach has proven beneficial in specific scenarios, such as games (1), (3) and (6). In the “(3) Divide the Dollar” game, incorporating CoT reduces the model’s propensity to suggest disproportionately large allocations. Similarly, in the “(6) Sealed-Bid Auction” game, CoT prompts the model to recognize bidding one’s true valuation as the optimal strategy. However, in the “(4) Public Goods Game” and “(5) Diner’s Dilemma,” the application of CoT inadvertently leads to a more self-serving model behavior, thereby undermining overall social welfare.
@@
Answer to RQ2: It is possible to improve gpt-3.5-0125 through simple prompt instructions. Among the methods we explore, the persona assignment of a “cooperative and collaborative assistant” performs the best.
@@ 5.4 RQ4: Leaderboard
This RQ investigates the variance in decision-making capabilities among different LLMs on $\gamma$-Bench. We analyze OpenAI’s GPT-3.5 across three iterations (i.e., 0613, 1106, and 0125), GPT-4 (0125), and Google’s Gemini Pro (1.0). The results are organized in Table 1, with model performance visualized in Fig. 14. Our findings reveal that GPT-4 markedly surpasses its counterparts, particularly in games (1), (3), (5), and (7). GPT-4’s diminished performance in the “(2) El Farol Bar” game stems from a conservative strategy favoring staying home. Its underperformance in the “(4) Public Goods Game” is attributed to a preference for individual gain over collective benefit. Furthermore, an evaluation of GPT-3.5’s iterations shows marked progress from version 0613 to versions 1106 and 0125, especially in the “(3) Divide the Dollar” and “(5) Diner’s Dilemma” games. Last but not least, we observe a discrepancy between Gemini Pro and GPT-4, primarily due to their performance in sequential games.
@@
Answer to RQ4: Currently, gpt-4-0125-preview outperforms all other models evaluated in this study, followed by gpt-3.5-turbo-1106. gemini-1.0-pro performs closely to gpt-3.5-turbo-0613.
@@ A.1 Cooperative Games
System
You are participating in a game played by $N$ players over $K$ rounds.
@@
User
Game Results for Round $I$:
@@
Assistant
{“decision”: “$D_{IJ}$”}
User
You gained $G_{IJ}$.
@@
User
Now round $I$ starts.
@@
System
You are participating in a game played by $N$ players over $K$ rounds.
@@
User
Game Results for Round $I$:
@@
Assistant
{“bid_amount”: “$B_{IJ}$”}
User
The sum of all bids was $S_{I}$.
@@
User
Now round $I$ starts.
@@ A.2 Betraying Games
System
You are participating in a game played by $N$ players over $K$ rounds.
@@
User
Game Results for Round $I$:
@@
Assistant
{“tokens_contributed”: “$C_{IJ}$”}
User
Tokens in the public pot: $S_{I}$
@@
User
Now round $I$ starts.
@@
System
You are participating in a game played by $N$ players over $K$ rounds.
@@
User
Game Results for Round $I$:
@@
Assistant
{“chosen_dish”: “$D_{IJ}$”}
User
Your utility is $u_{IJ}$.
@@
User
Now round $I$ starts.
@@
System
You are participating in a game played by $N$ players over $K$ rounds.
@@
User
Game Results for Round $I$:
@@
Assistant
{“bid”: “$b_{IJ}$”}
User
The winning bid was: $W_{I}$.
@@
User
Now round $I$ starts.
@@ A.3 Sequential Games
System
You are participating in a game played by $N$.
@@
User
Game Results for Round $I$:
@@
Assistant
(Only for the player itself) {“target”: “$t_{IJ}$”}
User
$NAME_{J}$ [intentionally missed the shot]/[shot at $t_{IJ}$ and hit]/[shot at $t_{IJ}$ but missed].
@@
User
Now round $I$ starts.
@@
System
You are participating in a game played by $N$.
@@
User
The $I$-th most senior pirate proposed a plan of {“$I$”: “$g_{II}$”, “$I+1$”: “$g_{II+1}$”, $\cdots$, “$I$”: “$g_{IN}$”}.
@@
Assistant
{“decision”: “$D_{I}J$”}
User
Less than half of the pirates accepted the plan.
@@
User
Now the $I$-th most senior pirate needs to propose a plan.
@@
The proposed plan is {“$I$”: “$g_{II}$”, “$I+1$”: “$g_{II+1}$”, $\cdots$, “$I$”: “$g_{IN}$”}. You will get $g_{IJ}$ golds from this plan.
@@
You need to propose a plan to divide $G$ golds. The proposed numbers must be all non-negative integers and sum up to $G$.

### v2

@@ Abstract
Decision-making, a complicated task requiring various types of abilities, presents an excellent framework for assessing Large Language Models (LLMs). Our research investigates LLMs’ decision-making capabilities through the lens of a well-established field, Game Theory. We focus specifically on games that support the participation of more than two agents simultaneously. Subsequently, we introduce our framework, $\gamma$-Bench, including eight classical multi-agent games. We design a scoring scheme to assess a model’s performance in these games quantitatively. Through $\gamma$-Bench, we investigate LLMs’ robustness, generalizability, and enhancement strategies. Results reveal that while GPT-3.5 shows satisfying robustness, its generalizability is relatively limited. However, its performance can be improved through approaches such as Chain-of-Thought. Additionally, we conduct evaluations across various LLMs and find that GPT-4 outperforms other models on $\gamma$-Bench, achieving a score of 60.5. Moreover, Gemini-1.0-Pro and GPT-3.5 (0613, 1106, 0125) demonstrate similar intelligence on $\gamma$-Bench. The code and experimental results are made publicly available via https://github.com/CUHK-ARISE/GAMABench.
@@ 1 Introduction
Public Goods Game $38.8_{\pm 8.1}$ $43.5_{\pm 12.6}$ $17.8_{\pm 1.7}$ $89.2_{\pm 1.8}$ $68.5_{\pm 7.6}$
Diner’s Dilemma $67.0_{\pm 4.9}$ $1.4_{\pm 1.3}$ $2.8_{\pm 2.8}$ $0.9_{\pm 0.7}$ $3.1_{\pm 1.5}$
Sealed-Bid Auction $5.5_{\pm 1.0}$ $3.4_{\pm 1.0}$ $4.2_{\pm 0.3}$ $8.7_{\pm 0.7}$ $14.9_{\pm 6.6}$
@@ 2.1 Game Theory
Capraro et al. (2023) ChatGPT-4, Bard, Bing Chat - Dictator Game
(Three Variants)
@@
Zhang et al. (2024) gpt-4 0.7 Guess 0.8 of the Average
Survival Auction Game
Duan et al. (2024) gpt-3.5-turbo, gpt-4, 0.2 Ten Gamesa
@@
Xie et al. (2024) text-davinci-003, - Seven Gamesb
gpt-3.5-turbo-instruct,
gpt-3.5-turbo-0613, gpt-4,
llama-2-(7/13/70)b,
vicuna-(7/13/33)b-v1.3
This Study gpt-3.5-turbo, gpt-4 0$\sim$1 Eight Gamesc
@@
  • a
    Tic-Tac-Toe, Connect-4, Kuhn Poker, Breakthrough, Liar’s Dice, Blind Auction, Negotiation, Nim, Pig, Iterated Prisoner’s Dilemma.
  • b
    Trust Game, Minimum Acceptable Probabilities Trust Game, Repeated Trust Game, Dictator Game, Risky Dictator Game, Lottery People Game, Lottery Gamble Game.
  • c
    Guess 2/3 of the Average, El Farol Bar, Divide the Dollar, Public Goods Game, Diner’s Dilemma, Sealed-Bid Auction, Battle Royale, Pirate Game.
Nash Equilibrium
Theorem 2.1 (Nash’s Existence Theorem)
Human Behaviors
@@ 3.2 Betraying Games
The Sealed-Bid Auction (SBA) involves players submitting their bids confidentially and simultaneously, different from the auctions where bids are made openly in a sequential manner. We consider two variants of SBA: the First-Price Sealed-Bid Auction (FPSBA) and the Second-Price Sealed-Bid Auction (SPSBA). In FPSBA, also known as the Blind Auction, if all players bid their true valuation $v_{i}$ of the item, the winner achieves a net gain of $b_{i}-v_{i}=0$ while others also gain nothing McAfee & McMillan (1987). Moreover, the highest bidder will discover that to win the auction, it is sufficient to bid marginally above the second-highest bid. Driven by these two factors, FPSBA is often deemed inefficient in practical scenarios, as bidders are inclined to submit bids significantly lower than their actual valuation, resulting in suboptimal social welfare. In contrast, SPSBA, commonly called the Vickrey auction, requires the winner to pay the second-highest bid, encouraging truthful bidding by all players Vickrey (1961). It can be proven that bidding true valuations in SPSBA represents an NE. This auction evaluates agent performance in imperfect information games, where agents lack knowledge of other players’ valuations.
@@ 4 Vanilla Experiments
Key Findings: The model’s decisions are mainly influenced by the outcomes of the preceding round rather than deriving from the reasoning of the optimal strategy. Although initially demonstrating suboptimal performance, the model can learn from historical data and enhance its performance over time. The model demonstrates the ability to engage in spontaneous cooperation, leading to increased social welfare beyond mere self-interest, without the necessity for explicit communication. However, this phenomenon also results in low performance in Betraying Games. The model shows limitations in sequential games characterized by more complicated rules. The aggregate score of the model on $\gamma$-Bench is $41.9$.
@@
System You are participating in a game played by $N$ players over $K$ rounds.
@@
User Game Results for Round $I$:
@@
Assistant {“chosen_number”: “$C_{IJ}$”}
User [Congratulation you won]/[Unfortunately you lost].
@@
User Now round $I$ starts.
@@ 4.2 Betraying Games
The vanilla setting for this game is $R=2$. Each player has $T=20$ to contribute in each round. Fig. 5 shows the contributed tokens by each agent and their corresponding gains per round. The observations reveal the following: (1) Despite an investment return of $-80\%$, agents display a pattern of alternating between free-riding and contributing all their tokens. (2) As the rounds progress, there is an evident increase in the number of tokens contributed to the public pot, leading to an overall enhancement in social welfare gains. These findings suggest that the LLM exhibits cooperative behavior, prioritizing collective benefits over individual self-interest. Since we expect the model to infer the optimal strategy, i.e., contributing zero tokens, the raw score in this game is given by $S_{4}=\frac{1}{NK}\sum_{ij}C_{ij}$, where $C_{ij}=1$ is the proposed contribution amount of player $i$ in round $j$. The model scores $41.3$ on this game.
@@
The vanilla setting for this game is $P_{h}=20$, $P_{l}=10$, $U_{h}=20$, $U_{l}=15$. We show the probability of agents choosing the costly dish, their resulting utilities, and the average bill in Figure 6. Analysis of the figure reveals the following insights: (1) Contrary to the NE predictions for this game, agents predominantly prefer the cheap dish, which maximizes total social welfare. (2) Remarkably, a deviation from cooperative behavior is observed wherein one agent consistently chooses to betray others, thereby securing a higher utility. This pattern of betrayal by this agent persists across subsequent rounds. Since we expect the model to infer the the optimal strategy, i.e., choosing the costly dish, the raw score in this game is given by $S_{5}=\frac{1}{NK}\sum_{ij}D_{ij}$, where $D_{ij}=1$ when player $i$ chose the cheap dish in round $j$ and $D_{ij}=0$ when player $i$ chose the costly dish. The model scores $4.0$ on this game.
@@
For the vanilla setting in this game, we randomly assign valuations to each agent in each round, ranging from 00 to $200$. We fix the seed for random number generation to ensure fair comparisons across various settings and models. We evaluate LLMs’ performance under both First-Price and Second-Price settings. Fig. 7 depicts the subtraction between valuations and bids and bid amounts of each agent. Our key findings include: (1) As introduced in §3.2, we note that agents generally submit bids that are lower than their valuations in the First-Price auction, a tendency indicated by the positive discrepancies between valuations and bids depicted in Fig. 7(a). (2) Though the NE suggests that everyone bids the amount of their valuation in the Second-Price setting, we find a propensity for bidding below valuation levels, as demonstrated in Fig. 7(c). Since the optimal strategy is to bid the prices lower than their true valuations [^4] We evaluate only the First-Price setting according to the definition of Betraying Games., the raw score in this game is given by $S_{6}=\frac{1}{NK}\sum_{ij}(v_{ij}-b_{ij})$, where $v_{ij}$ and $b_{ij}$ are player $i$’s valuation and bid in round $j$, respectively. The model scores $6.5$ on this game.
@@ 5.1 RQ1: Robustness
We also investigate the impact of prompt phrasing on model performance. Leveraging GPT-4, we rewrite our default prompt templates described in §A, generating four additional versions. We perform a manual checking process on the generated versions to ensure GPT-4’s adherence to game rules without altering critical data. The prompt templates can be found in §B in the appendix. We plot the results of using these templates in Fig. 11 and record the quantitative scores in Table 7 in the Appendix. Notably, our findings indicate that prompt wording can significantly affect performance, as shown by the performance fluctuations in Fig. 11(1), (5), and (6).
@@ 5.2 RQ2: Reasoning Strategies
According to Kojima et al. (2022), introducing a preliminary phrase, “Let’s think step by step,” encourages the model to sequentially analyze and explain its reasoning before presenting its conclusion. This approach has proven beneficial in specific scenarios, such as games (1), (3), (4), and (5). In the “(3) Divide the Dollar” game, incorporating CoT reduces the model’s propensity to suggest disproportionately large allocations. Similarly, in the “(4) Public Goods Game” and “(5) Diner’s Dilemma,” CoT prompts the model to recognize being a free-rider as the optimal strategy.
@@
Answer to RQ2: It is possible to improve gpt-3.5-0125 through simple prompt instructions. Among the methods we explore, the CoT prompting performs the best, achieving a performance close to GPT-4 ($57.4$ vs. $60.5$).
@@ 5.4 RQ4: Leaderboard
This RQ investigates the variance in decision-making capabilities among different LLMs on $\gamma$-Bench. We analyze OpenAI’s GPT-3.5 across three iterations (i.e., 0613, 1106, and 0125), GPT-4 (0125), and Google’s Gemini Pro (1.0). The results are organized in Table 1, with model performance visualized in Fig. 14. Our findings reveal that GPT-4 markedly surpasses its counterparts, particularly in games (1), (3), (4), (7) and (8). GPT-4’s diminished performance in the “(2) El Farol Bar” game stems from a conservative strategy favoring staying home. Its underperformance in the “(5) Diner’s Dilemma” game is attributed to a preference for individual gain over collective benefit. Furthermore, an evaluation of GPT-3.5’s iterations shows similar performance. Last but not least, we observe a discrepancy between Gemini Pro and GPT-4, primarily due to their performance in sequential games.
@@
Answer to RQ4: Currently, gpt-4-0125-preview outperforms all other models evaluated in this study. gemini-1.0-pro performs closely to all the GPT-3 versions.
@@ A.1 Cooperative Games
System You are participating in a game played by $N$ players over $K$ rounds.
@@
User Game Results for Round $I$:
@@
Assistant {“decision”: “$D_{IJ}$”}
User You gained $G_{IJ}$.
@@
User Now round $I$ starts.
@@
System You are participating in a game played by $N$ players over $K$ rounds.
@@
User Game Results for Round $I$:
@@
Assistant {“bid_amount”: “$B_{IJ}$”}
User The sum of all bids was $S_{I}$.
@@
User Now round $I$ starts.
@@ A.2 Betraying Games
System You are participating in a game played by $N$ players over $K$ rounds.
@@
User Game Results for Round $I$:
@@
Assistant {“tokens_contributed”: “$C_{IJ}$”}
User Tokens in the public pot: $S_{I}$
@@
User Now round $I$ starts.
@@
System You are participating in a game played by $N$ players over $K$ rounds.
@@
User Game Results for Round $I$:
@@
Assistant {“chosen_dish”: “$D_{IJ}$”}
User Your utility is $u_{IJ}$.
@@
User Now round $I$ starts.
@@
System You are participating in a game played by $N$ players over $K$ rounds.
@@
User Game Results for Round $I$:
@@
Assistant {“bid”: “$b_{IJ}$”}
User The winning bid was: $W_{I}$.
@@
User Now round $I$ starts.
@@ A.3 Sequential Games
System You are participating in a game played by $N$.
@@
User Game Results for Round $I$:
@@
Assistant (Only for the player itself) {“target”: “$t_{IJ}$”}
User $NAME_{J}$ [intentionally missed the shot]/[shot at $t_{IJ}$ and hit]/[shot at $t_{IJ}$ but missed].
@@
User Now round $I$ starts.
@@
System You are participating in a game played by $N$.
@@
User The $I$-th most senior pirate proposed a plan of {“$I$”: “$g_{II}$”, “$I+1$”: “$g_{II+1}$”, $\cdots$, “$I$”: “$g_{IN}$”}.
@@
Assistant {“decision”: “$D_{I}J$”}
User Less than half of the pirates accepted the plan.
@@
User Now the $I$-th most senior pirate needs to propose a plan.